A 5kg block and a 3kg block are connected by a string which hangs over a pulley.
a) How much weight does the 5 kg block have?
b) How much weight does the 3 kg block have?
c) How much weight is not balanced?
d) How much mass will be moved by this unbalanced force?
e) What acceleration will develop?
f) How much force acts in the string where it is tied to the 3 kg block?
g) How much force acts in the string where it is tied to the 5 kg block?
1
Expert's answer
2019-11-19T07:45:59-0500

(a) The weight of the 5 kg block


"W_1=m_1g=5\times 9.8=49\:\rm N"

(b) The weight of the 3 kg block


"W_2=m_2g=3\times 9.8=29.4\:\rm N"

(c) The not balanced weight


"\Delta W=W_1-W_2=49-29.4=19.6\:\rm N"

(d) The total mass


"m=m_1+m_2=5+3=8\:\rm kg"

(e) The acceleration


"a=\frac{\Delta W}{m}=\frac{19.6}{8}=2.45\:\rm m/s^2"

(f) The tension in the string 2


"T_2=m_2a+W_2=3\times 2.45+29.4=36.75\:\rm N"

(g) The tension in the string 1


"T_1=-m_1a+W_1=-5\times 2.45+49=36.75\:\rm N"